# Dynamic Post-Hoc Neural Ensemblers

### v1

@@ D.1 Finetuning Text Classifiers (FTC) Metadataset
It comprises three hyperparameters: model, learning rate and LoRA rank (Hu et al., 2022). We consider consider five models: 1) GPT2, 124M parameters; (Radford et al., 2019); 2) Bert-Large, 336M ; (Devlin et al., 2018); 3) Bart-Large 400 M, parameters  (Lewis et al., 2019); 4) Albert-Large, 17M parameters  (Lan et al., 2019); and 5) T5-Large, 770 M parameters  (Raffel et al., 2020). For the other two hyperparameters we also consider five different values as specified in Table 6.
Datasets
The metadataset contains predictions of models finetuned in five metadatasets for text classification: 1) IMDB (Maas et al., 2011); 2) Tweet (Maggie, 2020), 3) News (Zhang et al., 2015), 4) DBpedia (Zhang et al., 2015), 5) SST2 (Socher et al., 2013) and 6) SetFit (Tunstall et al., 2021). For every dataset, we create two versions. The first one is trained with the complete training data, while the second version is only with a subset of 10 % of the samples. All the datasets are for text classification beween 2 to 14 classes, including diverse domains such as movies, reviews, news, tweets, and text entailment data. We provide further information about the datasets in Table 7.
Figure 7: Mean error across datasets.
Metadataset Creation and Composition
We created the dataset by finetuning every model to the train split and, subsequently, saving their predictions on the validation and test split. The validation split corresponds to 20 % of the available train data. As for some datasets (SST-2 and SetFit) the test data is not completely provided by the creators, or it has hidden labels, we obtain it by using 20 % of the remaining test data. Specifically The models are fitted up to 5 epochs, using 1 GPU A100 with batch size equal to 2 and no LoRA dropout. We vary only the model, learning rate, and LoRA rank, while keeping the default hyperparameters in the Trainer object from the Transformers Library (version 4.41.0). Although we evaluate the models in a grid, some runs yielded out-of-memory errors for some configurations. In total, the metadataset contains 1134 evaluated configurations, representing around 3800 GPU hours. Additionally, we report information about the metadataset in Table 7.
Table 7: Metadataset Information
Dataset # Classes # Train Samples # Val. Samples # Test Samples # Confs (100%) # Confs. (10%)
IMDB (Maas et al., 2011) 2 20,000 5,000 25,000 125 125
Tweet (Maggie, 2020) 3 27,485 5,497 3,534 100 100
News (Zhang et al., 2015) 4 96,000 24,000 7,600 99 120
DBpedia (Zhang et al., 2015) 14 448,000 112,000 70,000 25 65
SST-2  (Socher et al., 2013) 2 43,103 13,470 10,776 125 125
SetFit  (Tunstall et al., 2021) 3 393,116 78,541 62,833 25 100
Hyperparameter Importance
We explore the importance of two hyperparameters, learning rate, and LoRA rank, by plotting the mean error as a heatmap in Figure 7. The error corresponds to the average across different models and datasets. The learning rate is an important hyperparameter, while increasing the LoRA rank does not affect the performance significantly in low learning rates. This behaviour is interesting, as it showcases that a small rank is enough for a successful finetuning in this context.
D.2 Scikit learn Pipelines
Search Space
@@ D.2 Scikit learn Pipelines
Our primary motivation is to investigate the ensembling of automated machine learning pipelines to enhance performance across various classification tasks. To effectively study ensembling methods and benchmark different strategies, we require a diverse set of pipelines. Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al., 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers. The pipelines are structured in three stages—preprocessor, feature selector, and classifier—which allows us to systematically explore numerous configurations. This extensive and diverse search space enables us to examine the impact of ensembling on a variety of models and serves as a robust benchmark for evaluating different ensembling techniques. Detailed descriptions of the components and their hyperparameters are provided in Tables 8, 9, and 10.
@@
Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al., 2019) as the foundation for our meta-dataset. OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains. This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
Metadataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space. Each pipeline execution was constrained to a maximum runtime of $15$ minutes. During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints. For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features. We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset. This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
Appendix E Additional Results
In this Section we report additional results from our experiments:
  • Average Ranking of baselines for the Negative Log-likelihood (Table 11) and Classification Errors 12.
  • Average NLL in Scikit-learn pipelines metadataset 8.
  • Average NLL using a subset of base models selected via DivBO 13, random 14 or both 15.
Table 11: Average Ranked NLL
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
Single-Best 14.0000±0.8944 14.0000±1.0000 12.8333±1.0408 6.8833±2.3219 7.3667±2.5049 7.5361±3.8531 7.7647±3.9096
Random 16.0000±0.0000 7.6667±4.0415 9.3333±0.5774 15.2500±0.8068 15.2333±0.7739 12.0000±3.9962 13.6176±1.8669
Top5 11.3333±2.5033 9.6667±1.1547 7.0000±1.0000 6.9000±1.9360 8.3000±1.9678 6.2470±3.3951 7.2353±4.2907
Top50 10.8333±1.7224 4.6667±1.1547 4.3333±0.5774 11.4000±1.2959 11.5333±1.5025 6.8554±3.4006 6.8824±3.7397
Quick 6.1667±1.6021 6.3333±3.0551 3.6667±2.5166 4.7333±2.0331 6.0333±2.5661 6.2048±2.8531 5.9412±3.2107
Greedy 3.5000±0.8367 4.0000±2.6458 12.8333±1.0408 3.5167±1.7786 4.5667±2.4238 5.9036±3.0268 6.5882±3.2607
CMAES 12.1667±4.4460 14.0000±1.0000 12.8333±1.0408 12.9667±2.6811 13.9667±1.3322 9.9277±3.8880 6.2353±2.4630
Random Forest 7.0000±2.7568 13.0000±2.6458 14.0000±2.6458 11.9833±1.2964 9.3000±3.8699 11.4217±5.1493 8.2353±4.8157
Gradient Boosting 3.1667±4.8339 14.6667±1.5275 13.5000±2.1794 9.7000±5.2203 10.1167±5.1890 11.4217±4.9463 8.3529±4.7558
SVM 10.6667±1.3663 10.0000±7.0000 11.3333±7.2342 13.9167±1.2532 11.6833±5.0198 10.8434±4.2641 11.7941±5.5707
Linear 7.6667±2.0656 11.0000±1.0000 10.0000±2.0000 6.8000±2.8816 6.5833±3.3479 10.3855±5.2587 14.3529±3.4989
MA 11.1667±2.8577 6.6667±1.1547 4.3333±2.3094 13.2167±1.0059 12.8167±1.6000 9.0241±4.0567 8.2941±5.1813
DivBO 7.1667±3.5449 9.0000±4.5826 5.3333±5.7735 4.9167±2.5123 5.1000±2.7114 6.6145±2.9377 6.7647±2.9054
EO 5.8333±3.7103 5.3333±1.5275 7.3333±1.5275 5.8833±2.2194 5.5000±2.5052 6.2169±2.6689 7.0000±2.5739
NE-Stack (Ours) 6.0000±4.7329 1.0000±0.0000 3.3333±4.0415 3.2000±3.4381 2.3333±2.0899 10.0241±5.8351 12.2941±3.5314
NE-MA (Ours) 3.3333±2.5033 5.0000±3.0000 4.0000±1.7321 4.7333±2.0331 5.5667±1.9945 5.3735±3.2674 4.6471±2.3702
Table 12: Average Ranked Error
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
Single-Best 12.1667±2.5626 15.0000±0.8660 13.8333±0.7638 9.8167±3.7908 9.1167±3.5591 8.9157±4.1480 9.3086±4.1030
Random 15.8333±0.4082 13.0000±0.0000 12.0000±1.0000 15.9667±0.1826 15.9667±0.1826 11.3253±5.0723 12.2716±4.3747
Top5 6.0833±4.2710 11.6667±2.0817 8.3333±2.0817 4.4000±2.8780 4.6167±2.9204 8.7048±3.6325 7.2099±4.0519
Top50 13.5000±2.2583 3.3333±1.5275 2.0000±1.0000 8.3000±4.1618 9.7333±2.6546 7.9880±4.1548 7.2222±3.6929
Quick 6.3333±4.1433 7.6667±1.1547 6.3333±2.3094 5.5833±3.5186 3.5500±2.3575 8.3253±3.1771 6.8148±3.8033
Greedy 5.0833±3.8264 11.3333±3.0551 13.8333±0.7638 9.0000±3.5012 6.9833±3.1961 8.0843±3.3467 6.9321±3.2984
CMAES 4.5833±2.2004 5.0000±3.6056 6.3333±2.0817 8.5833±3.1103 6.3167±2.1794 8.8313±3.7776 9.6605±4.5537
Random Forest 6.6667±4.4121 8.3333±4.6188 9.5000±4.4441 11.7000±3.6237 12.5000±1.9343 8.8313±4.0817 8.7778±4.5600
Gradient Boosting 10.5000±3.2711 15.6667±0.2887 14.5000±1.5000 8.4667±4.2729 12.4000±3.1139 9.1024±4.9840 9.0123±5.0812
SVM 3.5000±3.3317 8.0000±1.7321 8.6667±5.1316 10.1167±4.8382 13.3667±2.6682 7.0602±3.9296 10.5185±5.4906
Linear 7.8333±2.6394 11.6667±3.5119 11.3333±2.0817 5.7500±3.2450 7.5167±2.9870 7.7892±5.2179 9.0185±5.1705
MA 12.8333±2.4014 4.6667±0.5774 2.3333±1.5275 10.0667±3.5349 11.3667±2.9271 8.4277±4.2949 8.7531±3.7200
DivBO 9.4167±3.3229 10.6667±1.1547 12.6667±1.5275 10.0833±2.9916 7.3833±3.2262 8.1506±3.6599 8.4938±3.4228
EO 10.0000±4.6476 4.6667±3.2146 4.0000±2.0000 9.3667±2.8555 8.5000±2.6425 8.1988±4.2677 7.6852±2.8410
NE-Stack (Ours) 4.6667±3.0768 3.3333±1.1547 7.3333±0.5774 3.4500±2.3095 2.7167±2.2232 8.9639±5.4352 7.6173±5.5437
NE-MA (Ours) 7.0000±1.7889 2.0000±1.0000 3.0000±2.0000 5.3500±3.7374 3.9667±2.2967 7.3012±3.9900 6.7037±3.7855
Table 13: Average NLL for Subset of Base Models with DivBO
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
Table 14: Average NLL for Subset of Base Models randomly
Seletor FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
Table 15: Average NLL for Subset of Base Models (more baselines)
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single DivBO 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9767±0.0404 0.8220±0.3507 0.8814±0.2062 0.9748±0.1415 1.0287±0.1365
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9805±0.0339 0.7916±0.3245 0.8569±0.2178 0.9672±0.1381 1.0232±0.1550
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9687±0.0542 0.8048±0.3271 0.8672±0.2195 0.9652±0.1353 1.0378±0.1959
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9805±0.0339 0.7910±0.3209 0.8572±0.2174 0.9673±0.1384 1.0256±0.1601
CMAES DivBO 0.8915±0.1759 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 0.9671±0.1179 1.0250±0.1595
Random Forest DivBO 0.7932±0.1194 0.9338±0.3436 1.1678±0.2906 1.6807±1.1160 1.5507±0.8449 1.3195±0.6428 1.0534±0.1277
Gradient Boosting DivBO 0.7908±0.1848 1.4011±0.5359 1.0000±0.0000 2.4553±1.1191 2.7220±1.3007 1.2106±0.5937 1.1246±0.4052
Linear DivBO 0.7433±0.0870 0.6471±0.2272 1.0248±0.0430 0.9690±0.4209 1.1132±0.4525 1.0814±1.0056 1.0316±0.1403
SVM DivBO 0.8312±0.0943 0.7406±0.2612 1.0786±0.1362 5.2463±3.4452 4.8194±3.5169 1.5156±1.5597 2.7777±2.9458
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9719±0.0486 0.7603±0.2876 0.8802±0.1741 0.9621±0.1306 1.0222±0.1570
Single Random 1.0067±0.0164 1.0000±0.0000 1.0540±0.0936 0.8958±0.3297 0.9786±0.2444 0.9902±0.1529 1.1314±0.4475
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.7283±0.1534 2.7131±1.5510 2.5961±1.9153 1.2873±1.5773 1.0588±0.2762
Quick Random 0.7305±0.0764 0.5958±0.1917 0.7467±0.1420 0.9847±0.1836 0.9737±0.1340 1.0043±0.2497 1.0123±0.1529
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.7478±0.1352 0.9171±0.0932 0.8960±0.0795 0.9762±0.1082 1.0002±0.1038
CMAES Random 1.2164±0.3851 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.1544±0.9890 0.9985±0.0954
RF Random 0.7505±0.0915 0.8325±0.2255 0.9850±0.0175 2.9838±1.7970 2.6001±1.8694 1.2006±0.4367 1.0058±0.0898
GBT Random 0.7235±0.1605 1.8377±1.4510 1.2177±0.3771 1.9521±1.2076 2.8683±1.9763 1.2329±0.6934 1.0470±0.1913
Linear Random 0.7541±0.0897 0.7400±0.2827 0.9103±0.1384 1.1776±0.3187 1.5498±0.8446 1.0417±1.0146 1.0408±0.1296
SVM Random 0.8010±0.0903 0.7767±0.3006 0.9821±0.3676 5.1993±3.4269 5.1080±3.2902 1.4479±1.4990 2.8328±3.0214
MA Random 0.9069±0.1812 0.8677±0.2292 0.7543±0.1312 4.1694±2.5844 3.0315±2.4578 1.5059±1.9654 1.4185±1.6833
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6882±0.1602 0.6925±0.4510 0.6854±0.2801 0.9387±0.4628 4.2065±3.4678
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.7535±0.1164 0.9291±0.0999 0.9191±0.0624 0.9922±0.2525 1.0259±0.2002
NE-Stack DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.0442±0.8691 0.7482±0.2254 0.9671±0.3746 6.9574±3.4368
NE-MA DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 0.7632±0.3319 0.8349±0.2075 0.9677±0.1455 1.0057±0.1121

### v2

@@ Appendix I Proof-of-Concept with Overparameterized Base Models
Our primary motivation is to investigate the ensembling of automated machine learning pipelines to enhance performance across various classification tasks. To effectively study ensembling methods and benchmark different strategies, we require a diverse set of pipelines. Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al.,, 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers. The pipelines are structured in three stages: preprocessor, feature selector, and classifier. This allows us to explore numerous configurations systematically. This extensive and diverse search space enables us to examine the impact of ensembling on a variety of models and serves as a robust benchmark for evaluating different ensembling techniques. Detailed descriptions of the components and their hyperparameters are provided in Tables 8, 9, and 10.
Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al.,, 2019) as the foundation for our meta-dataset. OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains. This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
Meta-Dataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space. Each pipeline execution was constrained to a maximum runtime of $15$ minutes. During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints. For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features. We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset. This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
@@
Table 11: Metadatasets Information
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset Avg. Classes per Dataset
Nasbench (100) Vision NAS, Classification (Dong and Yang,, 2020) 3 11000 6000 100 76.6
Nasbench (1K) Vision NAS, Classification (Dong and Yang,, 2020) 3 11000 6000 1K 76.6
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024a, ) 30 160 160 255 20.
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024a, ) 30 1088 1088 203 136.
FTC Language Finetuning, Classification, Section (Arango et al., 2024b, ) 6 39751 29957 105 4.6
TabRepo Clas. Tabular Classification  (Salinas and Erickson,, 2023) 83 1134 126 1530 3.4
TabRepo Reg. Tabular Regression  (Salinas and Erickson,, 2023) 17 3054 3397 1530 -
Sk-Learn Pipelines. Tabular Classification, Section O 69 1514 1514 500 5.08